Solve logarithmic equation. Log8x^3=4
THERE IS NO BASE. The equation is just like it is seen

Answer by Alan3354(69443)

Log8x^3=4
----------
Log(8x^3) = 4
8x^3 = 10^4 = 10000
x^3 = 1250
x =~ 10.7722